The current in an RL circuit decreases from 0.50 A to 5.0 mA in the one second after the battery is switched out of the circuit. If L = 11.4 H, what is the resistance in the circuit? Express your answer in ohms.
Answer #1

let T is the time constant of the circuit.

use,

I(t) = I_max*e^(-t/T)

5*10^-3 = 0.5*e^(-1/T)


10^-2 = e^(-t/T)

ln(10^-2) = -t/T

==> T = -t/ln(10^-2)

= -1/ln(10^-2)

= 0.2171 s

we know, T = L/R

==> R = L/T

= 11.4/0.2171

= 52.5 ohms <<<<<<<<<<<<-----------------------Answer